In a coin collection there are 9 pennies, 8 nickels, 12 dimes, 16 quarters, 11 half-dollars and 4 dollar pieces. A coin is randomly selected from the collection by a blind folded boy who gets to keep the coin picked. What is the expected value?
Answer by ilana(307)

The expected value is the sum of each probability times its value.
P(penny) = 9/60
P(nickel) = 8/60
P(dime) = 12/60
P(quarter) = 16/60
P(half-dollar) = 11/60
P(dollar) = 4/60
Expected value = 1*9/60 + 5*8/60 + 10*12/60 + 25*16/60 + 50*11/60 + 100*4/60
= (9 + 40 + 120 + 400 + 550 + 400)/60
= 1519/60
= approx. $25.32